# Directed Exploration in Reinforcement Learning from Linear Temporal Logic

### v1

@@ 1 Introduction
Most reinforcement learning (RL) research has traditionally focused on a standard setting, prescribing the maximization of cumulative rewards in a Markov Decision Process [32, 40]. While this simple formalism captures a variety of behaviors [39], its expressiveness remains limited [1]. In pursuit of a more natural and effective way to specify desired behavior, several works have turned towards logic languages [28, 17, 7, 9, 22]. Originally designed to describe possible paths of a system (with direct applications, e.g., in model checking [3]), Linear Temporal Logic (LTL) [31] has been found to strike an interesting balance between expressiveness and tractability.
A translation from LTL to RL objectives is in general possible at the cost of optimality [2]. Nevertheless, several works [4, 18, 45] have proposed a reward and discounting scheme to distill a policy through RL from an LTL specification. In particular, the scheme proposed by Voloshin et al. [45] recovers a policy that optimizes a lower bound on the probability of satisfying the specification: suboptimality can in practice be bounded through discretization arguments, of for finite policy classes. However, this scheme results in a sparse reward signal and a potentially flat value landscape, thus making exploration a fundamental challenge.
The necessity for strong exploration algorithms when learning from LTL specification is therefore evident. Existing methods rely on counterfactual data augmentation schemes [45], which however do not directly guide the agent in the underlying MDP, on the availability of a task distribution [47], or on heuristics [18]. Closer in spirit to our work, task-aware reward shaping has briefly been explored in the context of finite logic [7, 22], but has not been scaled to $\omega$-regular problems and cannot adapt to unknown environment dynamics.
@@
Our method, named DRL2 (Directed Reinforcement Learning from Linear Temporal Logic), is illustrated in Figure 1 and is capable of driving deep exploration in reinforcement learning from linear temporal logic specifications. The contributions of this work can be outlined as follows:
@@
    we evaluate the method across diverse settings, spanning from simple tabular cases to complex, high-dimensional environments, which pose a significant challenge for RL from LTL specifications.
Section 2 provides an introduction to LTL and its connection to RL. Our method is described in Section 3 and evaluated in Section 4. A discussion of related works and of the proposed one can be found in Sections 5 and 6, respectively.
@@ 2 Background
This section provides a brief introduction to linear temporal logic and discusses connections to reinforcement learning. For a complete introduction, we refer the reader to Baier and Katoen [3].
@@
Each LTL formula can be satisfied by an infinite sequence of truth evaluations of AP (i.e., an $\omega$-word). While a direct definition is also possible [41], for simplicity, satisfaction will be introduced through the acceptance of paths in an automaton built from the formula.
@@
A practical way of defining satisfaction involves the introduction of Limit Deterministic Büchi Automata (LDBAs). An LDBA can be constructed from any LTL formula [38] and is able to keep track of the progression of its satisfaction.
@@
In LDBAs synthesized from LTL formulas, additional properties hold [38]. First, the alphabet is over evaluations of atomic propositions $\Sigma=2^{AP}$. Second, the non-determinism is limited to a subset of so-called $\epsilon$-transitions, which arise for specific formulas, e.g. those encoding stability problems. As a consequence, this class of LDBAs is known as Good-for-MDPs [15], since the non-determinism can be resolved on the fly in arbitrary MDPs without changing acceptance.
@@
We note that this notion of success relies on conditions that are achieved eventually in the future, independently of temporal distance. While this leads to myopic behavior under naive optimization [45], recent works [4, 18, 8, 45] propose an elegant rephrasing for formula satisfaction in an RL-friendly form, as we now describe.
@@
The following part describes standard RL terminology and then reconciles it with the introduced logic machinery.
@@
MDPs are the standard modeling choice for RL environments; however, they are disconnected from atomic propositions and unable to track progression over formula satisfaction. The semantics of APs can be grounded in MDP states through a labeling function $\mathcal{F}:\mathcal{S}\to\Sigma$, which evaluates APs in each MDP state. On the other hand, progression over task satisfaction can naturally be stored in an LDBA, as its states encode sufficient information on the history of paths.
Finally, all three components (MDP, labeling function and LDBA) can be synchronized to ensure consistency between trajectories in each of them and enable mapping policies to distribution over paths, and therefore likelihoods of formula satisfaction [17, 45]. The first step is to resolve the non-determinism by creating a new action for each potential $\epsilon$-transition to a state $b\in\mathcal{B}$, thus defining a new action set $\mathcal{A^{\mathcal{B}}}=\{\epsilon_{b}|b\in\mathcal{B}\}$ [38]. Then, we can define a product MDP ×$\mathcal{M^{\times}}=(\mathcal{S^{\times}},\mathcal{A^{\times}},\mathcal{P^{\times}},R^{\times},\gamma^{\times},\mu^{\times}_{0})$, where $\mathcal{S^{\times}}=\mathcal{S}\times\mathcal{B},\mathcal{A^{\times}}=\mathcal{A}\cup\mathcal{A^{B}}$, and $\mu^{\times}_{0}(s,b)=\mu_{0}(s)\cdot\mathbf{1}_{b=b_{0}}$.
Essentially, a policy learned over the product MDP has access to both low-level information (MDP states) and indicators of progress on the specified task (LDBA states). The transition kernel over $\mathcal{M^{\times}}$ needs to guarantee that both the underlying MDP and the LDBA evolve consistently. This synchronization is achieved through the labeling function $\mathcal{F}$:
@@
where $b\stackrel{{\scriptstyle\epsilon}}{{\to}}b^{\prime}$ indicates that there is an $\epsilon$-transition between the two LDBA states $b$ and $b^{\prime}$. Through this construction, it is finally possible to connect trajectories (and, therefore, policies) to satisfaction of a given LTL formula. Let us consider an LTL formula $\varphi$ and its corresponding LDBA $\mathcal{L}$, as well as a trajectory $\tau=(s_{i},b_{i})_{0}^{\infty}$ in the product MDP $\mathcal{M}^{\times}$. Then, $\tau\models\varphi$ ($\tau$ satisfies $\varphi$) if and only if $\mathcal{L}$ accepts the path $(b_{i})_{0}^{\infty}$, i.e., the projection of $\tau$ to LDBA states. Finally, let us consider a policy $\pi:\mathcal{S^{\times}}\to\Delta(\mathcal{A^{\times}})$: the probability of $\pi$ satisfying $\varphi$ can thus be defined as the probability integral for trajectories satisfying the formula: $P(\pi\models\varphi)=\mathop{\mathbb{E}}_{\tau\sim\pi}\mathbf{1}_{\tau\models\varphi}$. Optimizing a policy $\pi$ for satisfaction of an LTL specification $\varphi$ can be expressed as finding $\pi^{\star}\in\mathop{\text{argmax}}_{\pi\in\Pi}P(\pi\models\varphi)$ Prior works [17, 45] have proposed RL-friendly proxy objectives, which optimize a lower bound on the probability of LTL formula satisfaction:
$\pi^{\star}_{\gamma}\in\mathop{\text{argmax}}_{\pi\in\Pi}\mathop{\mathbb{E}}_{\tau\sim\pi}\Big{[}\sum_{i=0}^{\infty}\Gamma_{i}R^{\times}(b_{i})\Big{]}\;(:=V^{\gamma}_{\pi}),\text{ where}$ (1)
$R^{\times}(b_{i})=\mathbf{1}_{\{b_{i}\in\mathcal{B}^{*}\}},\;\Gamma_{0}=1,\;\Gamma_{i}=\prod_{t=0}^{i-1}\gamma^{\times}(b_{t}),\quad\gamma^{\times}(b_{t})=\begin{cases}\gamma,&b_{t}\in\mathcal{B}^{*}\\ 1,&\text{otherwise.}\end{cases}$ (2)
Paraphrasing, $\pi^{\star}_{\gamma}$ maximizes the visitation count to accepting states in the LDBA under eventual discounting. For a formal analysis of the policy recovered by this objective, we refer the reader to Voloshin et al. [45]. Our work builds upon this formulation and devises an exploration method to compensate for its drawbacks. That is, the reward function $R^{\times}$ is fundamentally sparse: the agent only receives feedback when a significant amount of progress toward solving the task has been made, and thus an accepting LDBA state is visited. As a result, while naive exploration might reach several non-accepting LDBA states, the agent remains unable to evaluate them, as it is largely uninformed of the yet unexplored parts of the LDBA. Our method distills the global known structure of the LDBA in a denser intrinsic reward signal for exploration.
@@ 3.1 High-level Value Estimation from LDBA
As described in Section 2, an LDBA can be naturally synthesized from a given LTL specification, through well-known schemes [38]. We now show how an LDBA can be recast as a Markov reward process, which can be leveraged for computing values for each LDBA state. This construction requires 2 ingredients, namely (i) an LDBA $\mathcal{L}$, and (ii) a transition kernel $K$ over LDBA states, where $K$ is a stochastic matrix such that $K_{i,j}=\mathop{\mathbb{E}}_{\pi,P^{\times},\mu^{\times}_{0}}P(b^{\prime}=b_{j}|b=b_{i})$, i.e., an estimate of the probability for the LDBA to transition to state $b_{j}$ starting from $b_{i}$ under some policy $\pi$, assuming stationarity. The choice of the transition kernel $K$ is crucial for the effectiveness of the method and is thus treated in detail in the following section.
@@
where a matrix notation is adopted: $\bar{V}_{K},R^{\times}$ and $\gamma^{\times}$ are represented as $|\mathcal{B}|$-dimensional vectors, and $\odot$ stands for the Hadamard product. As the LDBA (and thus the MRP) is discrete and finite, the Bellman equation defined over the MRP has a closed-form solution [40]  [^2] In the case of exceedingly complex specifications, and thus large LDBAs, iterative methods could be a viable replacement.:
@@
where $\mathbb{I}$ represents the identity matrix. As the MRP is closely related to the product MDP $\mathcal{M}^{\times}$, an analysis of their connection is provided in Appendix B.
Once value estimates $\bar{V}_{K}$ are computed, they can be treated as a potential function for reward shaping [29], although under eventual discounting [45]:
@@
This reward signal can be added to the product MDP reward $R^{\times}$ and optimized with an arbitrary RL algorithm. We note that, as an instantiation of potential-based reward shaping, the optimal policy in the product MDP remains invariant to this reward transformation, if eventual discounting converges to zero.
@@
The proof follows the general scheme for potential-based reward shaping [29] and extends it to the eventual discounting setting (see Appendix A).
@@
Figure 3: Top: a randomly initialized policy is executed in the product MDP $\mathcal{M}^{\times}$ for T0; its empirical transition kernel $\hat{K}$ results in uniformly zero value estimates $\bar{V}_{K}$. Bottom: the expected value for $K$ over a posterior distribution estimated from a symmetric prior results in informative value estimates and intrinsic rewards.
@@ 3.2 Optimistic Priors for High Level Value Estimation
In order to address this issue, we adopt a Bayesian approach to the problem of estimating the LDBA transition kernel $K$. Let us consider each row $K_{i}$, which models a categorical probability distribution over future LDBA states from each LDBA state $b_{i}\in\mathcal{B}$. At its core, our method proposes a prior distribution over these categoricals, such that appropriate shaping of the prior controls and directs the exploration in the product MDP. As the conjugate prior to categorical distributions, we adopt a Dirichlet prior $K_{i}\sim P(K_{i})=\textbf{Dir}(\alpha_{i,0},\dots,\alpha_{i,|\mathcal{B}|-1})$. The prior is informed of the LDBA structure by setting $\alpha_{i,j}=0$ if the LDBA does not allow transitions from $b_{i}$ to $b_{j}$. For the $m$ remaining non-zero Dirichlet parameters, we adopt an partially symmetric prior by setting them to a shared value $\frac{\alpha}{m}$, where $\alpha$ is a hyperparameter controlling the strength of the prior: large values of $\alpha$ induce slower convergence of the posterior distribution to the empirical transition kernel.
We remark that the choice of symmetry corresponds to the assumption that, at each step, the agent is capable of transitioning to each adjacent LDBA state with equal probability. In practice, the agent might actually take several steps in the underlying MDP in order to transition to any different LDBA state; moreover, some LDBA transitions can be substantially harder to achieve than others. Finally, for complex problems, naive exploration may not even result in observing the full set of possible transitions in a practical number of time steps. However, assuming that the agent is capable of transitioning under a max-entropy distribution allows reward signals to propagate to all states, thus resulting in informative estimates for the high-level value $\bar{V}_{K}$ and the intrinsic rewards $R_{\text{intr}}$. For the illustrative example $T_{0}$, this is displayed in the bottom part of Figure 3.
@@
for each $i\in[0,\dots,|\mathcal{B}|]$. This update is tractable due to the choice of a conjugate Dirichlet prior for the categorical likelihood described by the transition kernel. As training progresses, a posterior distribution $P(K|D)$ can be updated with collected evidence. Moreover, instead of computing high-level values for a specific transition kernel $K$ as in Equation 4, we can compute the expected value $\bar{V}$ under the posterior distribution of transition kernels $P(K|D)$ and thus of MRPs:
@@
We remark that the maximization of $\bar{V}$ corresponds to the average-case MDP problem, while other procedures can be easily adapted to solve the Robust MDP [50] or the Percentile Criterion MDP problems [11]. In practice, the expectation in Equation 7 can be estimated by sampling, with the special case of Thompson Sampling when a single sample is used. Our approach to estimating the transition kernel is ablated empirically in Appendix F; a study of the hyperparameter $\alpha$ controlling prior strength is in Appendix G.

### v2

@@ 1 Introduction
Most reinforcement learning (RL) research has traditionally focused on a standard setting, prescribing the maximization of cumulative rewards in a Markov Decision Process (Puterman, 2014, Sutton & Barto, 2018). While this simple formalism captures a variety of behaviors (Silver et al., 2021), its expressiveness remains limited (Abel et al., 2021). In pursuit of a more natural and effective way to specify desired behavior, several works have turned towards logic languages (Li et al., 2017, Hasanbeig et al., 2018, Camacho et al., 2019, De Giacomo et al., 2020b, Icarte et al., 2022). Originally designed to describe possible paths of a system (with direct applications, e.g., in model checking (Baier & Katoen, 2008)), Linear Temporal Logic (LTL) (Pnueli, 1977) has been found to strike an interesting balance between expressiveness and tractability.
A translation from LTL to reinforcement learning objectives is in general possible at the cost of optimality (Alur et al., 2022). Nevertheless, several works (Bozkurt et al., 2020, Hasanbeig et al., 2020, Voloshin et al., 2023) have proposed a reward and discounting scheme to distill a policy through reinforcement learning from an LTL specification. In particular, the scheme proposed by Voloshin et al. (2023) recovers a policy that optimizes a lower bound on the probability of satisfying the specification: suboptimality can in practice be bounded through discretization arguments, or for finite policy classes. However, this scheme results in a sparse reward signal and a potentially flat value landscape, thus making exploration a fundamental challenge.
The necessity for strong exploration algorithms when learning from LTL specification is therefore evident. Existing methods rely on counterfactual data augmentation schemes (Voloshin et al., 2023), which however do not directly guide the agent in the underlying MDP, on the availability of a task distribution (Wang et al., 2023), or on heuristics (Hasanbeig et al., 2020). Closer in spirit to our work, task-aware reward shaping has briefly been explored in the context of finite logic (Camacho et al., 2019, Icarte et al., 2022), but has not been scaled to $\omega$-regular problems and cannot adapt to unknown environment dynamics.
@@
Our method, named DRL2 (Directed Reinforcement Learning from Linear Temporal Logic), is illustrated in Figure 1 and is capable of driving deep exploration in reinforcement learning from LTL specifications. The contributions of this work can be outlined as follows:
@@
    we evaluate the method across diverse settings, spanning from simple tabular cases to complex, high-dimensional environments, which pose a significant challenge for reinforcement learning from LTL specifications.
Section 2 provides an introduction to LTL and its connection to reinforcement learning. Our method is described in Section 3 and evaluated in Section 4. A discussion of related works and of the proposed one can be found in Sections 5 and 6, respectively.
@@ 2 Background
This section provides a brief introduction to LTL and discusses connections to reinforcement learning. For a complete introduction, we refer the reader to Baier & Katoen (2008).
@@
Each LTL formula can be satisfied by an infinite sequence of truth evaluations of AP (i.e., an $\omega$-word). While a direct definition is also possible (Thomas, 1990), for simplicity, satisfaction will be introduced through the acceptance of paths in an automaton built from the formula.
@@
A practical way of defining satisfaction involves the introduction of Limit Deterministic Büchi Automata (LDBAs). An LDBA can be constructed from any LTL formula (Sickert et al., 2016) and is able to keep track of the progression of its satisfaction.
@@
In LDBAs synthesized from LTL formulas, additional properties hold (Sickert et al., 2016). First, the alphabet is over evaluations of atomic propositions $\Sigma=2^{AP}$. Second, the non-determinism is limited to a subset of so-called $\epsilon$-transitions, which arise for specific formulas, e.g. those encoding stability problems. As a consequence, this class of LDBAs is known as Good-for-MDPs (Hahn et al., 2020), since the non-determinism can be resolved on the fly in arbitrary MDPs without changing acceptance.
@@
We note that this notion of success relies on conditions that are achieved eventually in the future, independently of temporal distance. While this leads to myopic behavior under naive optimization (Voloshin et al., 2023), recent works (Bozkurt et al., 2020, Hasanbeig et al., 2020, De Giacomo et al., 2020a, Voloshin et al., 2023) propose an elegant rephrasing for formula satisfaction in an RL-friendly form, as we now describe.
@@
The following part describes standard reinforcement learning terminology and then reconciles it with the introduced logic machinery.
@@
MDPs are the standard modeling choice for reinforcement learning environments; however, they are disconnected from atomic propositions and unable to track progression over formula satisfaction. The semantics of APs can be grounded in MDP states through a labeling function $\mathcal{F}:\mathcal{S}\to\Sigma$, which evaluates APs in each MDP state. On the other hand, progression over task satisfaction can naturally be stored in an LDBA, as its states encode sufficient information on the history of paths.
Finally, all three components (MDP, labeling function and LDBA) can be synchronized to ensure consistency between trajectories in each of them and enable mapping policies to distribution over paths, and therefore likelihoods of formula satisfaction (Hasanbeig et al., 2018, Voloshin et al., 2023). The first step is to resolve the non-determinism by creating a new action for each potential $\epsilon$-transition to a state $b\in\mathcal{B}$, thus defining a new action set $\mathcal{A^{\mathcal{B}}}=\{\epsilon_{b}|b\in\mathcal{B}\}$ (Sickert et al., 2016). Then, we can define a product MDP, encompassing the underlying MDP and the LDBA.
Definition 2.5.
(Product MDP) Given an MDP $\mathcal{M}$ and a LDBA $\mathcal{L}$, their Product MDP is a tuple ×$\mathcal{M^{\times}}=(\mathcal{S^{\times}},\mathcal{A^{\times}},\mathcal{P^{\times}},R^{\times},\gamma^{\times},\mu^{\times}_{0})$, where $\mathcal{S^{\times}}=\mathcal{S}\times\mathcal{B},\mathcal{A^{\times}}=\mathcal{A}\cup\mathcal{A^{B}}$, and $\mu^{\times}_{0}(s,b)=\mu_{0}(s)\cdot\mathbf{1}_{b=b_{0}}$. Furthermore,
@@
where $b\stackrel{{\scriptstyle\epsilon}}{{\to}}b^{\prime}$ indicates that there is an $\epsilon$-transition between the two LDBA states $b$ and $b^{\prime}$. Finally, the reward and discount factor are defined as
$R^{\times}(b_{i})=\begin{cases}1,&\{b_{i}\in\mathcal{B}^{*}\}\\ 0,&\text{otherwise}\end{cases},\quad\gamma^{\times}(b_{t})=\begin{cases}\gamma,&b_{t}\in\mathcal{B}^{*}\\ 1,&\text{otherwise.}\end{cases}$ (1)
Essentially, a policy learned over the product MDP has access to both low-level information (MDP states) and indicators of progress on the specified task (LDBA states). The transition kernel over $\mathcal{M^{\times}}$ guarantees that both the underlying MDP and the LDBA evolve consistently. This synchronization is achieved through the labeling function $\mathcal{F}$. Through this construction, it is finally possible to connect trajectories (and, therefore, policies) to satisfaction of a given LTL formula. Let us consider an LTL formula $\varphi$ and its corresponding LDBA $\mathcal{L}$, as well as a trajectory $\tau=(s_{i},b_{i})_{0}^{\infty}$ in the product MDP $\mathcal{M}^{\times}$. Then, $\tau\models\varphi$ ($\tau$ satisfies $\varphi$) if and only if $\mathcal{L}$ accepts the path $(b_{i})_{0}^{\infty}$, i.e., the projection of $\tau$ to LDBA states. Finally, let us consider a policy $\pi:\mathcal{S^{\times}}\to\Delta(\mathcal{A^{\times}})$: the probability of $\pi$ satisfying $\varphi$ can thus be defined as the probability integral for trajectories satisfying the formula: $P(\pi\models\varphi)=\mathop{\mathbb{E}}_{\tau\sim\pi}\mathbf{1}_{\tau\models\varphi}$. Optimizing a policy $\pi$ for satisfaction of an LTL specification $\varphi$ can be expressed as finding $\pi^{\star}\in\mathop{\text{argmax}}_{\pi\in\Pi}P(\pi\models\varphi)$ Prior works (Hasanbeig et al., 2018, Voloshin et al., 2023) have proposed RL-friendly proxy objectives, which optimize a lower bound on the probability of LTL formula satisfaction:
$\pi^{\star}_{\gamma}\in\mathop{\text{argmax}}_{\pi\in\Pi}\mathop{\mathbb{E}}_{\tau\sim\pi}\Big{[}\sum_{i=0}^{\infty}\Gamma_{i}R^{\times}(b_{i})\Big{]}\;(:=V^{\gamma}_{\pi}),\quad\text{where}\;\Gamma_{0}=1,\;\Gamma_{i}=\prod_{t=0}^{i-1}\gamma^{\times}(b_{t}).$ (2)
Paraphrasing, $\pi^{\star}_{\gamma}$ maximizes the visitation count to accepting states in the LDBA under eventual discounting. For a formal analysis of the policy recovered by this objective, we refer the reader to Voloshin et al. (2023). Our work builds upon this formulation and devises an exploration method to compensate for its drawbacks. That is, the reward function $R^{\times}$ is fundamentally sparse: the agent only receives feedback when a significant amount of progress toward solving the task has been made, and thus an accepting LDBA state is visited. As a result, while naive exploration might reach several non-accepting LDBA states, the agent remains unable to evaluate them, as it is largely uninformed of the yet unexplored parts of the LDBA. Our method distills the global known structure of the LDBA in a denser intrinsic reward signal for exploration.
@@ 3.1 High-level Value Estimation from LDBA
As described in Section 2, an LDBA can be naturally synthesized from a given LTL specification, through well-known schemes (Sickert et al., 2016). We now show how an LDBA can be recast as a Markov reward process, which can be leveraged for computing values for each LDBA state. This construction requires 2 ingredients, namely (i) an LDBA $\mathcal{L}$, and (ii) a transition kernel $K$ over LDBA states, where $K$ is a stochastic matrix such that $K_{i,j}=\mathop{\mathbb{E}}_{\pi,P^{\times},\mu^{\times}_{0}}P(b^{\prime}=b_{j}|b=b_{i})$, i.e., an estimate of the probability for the LDBA to transition to state $b_{j}$ starting from $b_{i}$ under some policy $\pi$, assuming stationarity. The choice of the transition kernel $K$ is crucial for the effectiveness of the method and is thus treated in detail in the following section.
@@
where a matrix notation is adopted: $\bar{V}_{K},R^{\times}$ and $\gamma^{\times}$ are represented as $|\mathcal{B}|$-dimensional vectors, and $\odot$ stands for the Hadamard product. As the LDBA (and thus the MRP) is discrete and finite  [^2] We remark that the MRP is constructed from the LDBA, which is discrete in nature, and not from the underlying MDP, which can be potentially continuous., the Bellman equation defined over the MRP has a closed-form solution (Sutton & Barto, 2018)  [^3] In the case of exceedingly complex specifications, and thus large LDBAs, iterative methods could be a viable replacement.:
@@
where $\mathbb{I}$ represents the identity matrix. As the MRP is closely related to the product MDP $\mathcal{M}^{\times}$, an analysis of their connection is provided in Appendix B. Once value estimates $\bar{V}_{K}$ are computed, they can be treated as a potential function for reward shaping (Ng et al., 1999), although under eventual discounting (Voloshin et al., 2023):
@@
This reward signal can be added to the product MDP reward $R^{\times}$ and optimized with an arbitrary reinforcement learning algorithm. We note that, as an instantiation of potential-based reward shaping, the optimal policy in the product MDP remains invariant to this reward transformation, if eventual discounting converges to zero.
@@
The proof follows the general scheme for potential-based reward shaping (Ng et al., 1999) and extends it to the eventual discounting setting (see Appendix A).
@@
Figure 3: Left: a randomly initialized policy is executed in the product MDP $\mathcal{M}^{\times}$ for T0; its empirical transition kernel $\hat{K}$ results in uniformly zero value estimates $\bar{V}_{K}$. Right: the expected value for $K$ over a posterior distribution estimated from a symmetric prior results in informative value estimates and intrinsic rewards.
@@ 3.2 Optimistic Priors for High Level Value Estimation
In order to address this issue, we adopt a Bayesian approach to the problem of estimating the LDBA transition kernel $K$. Let us consider each row $K_{i}$, which models a categorical probability distribution over future LDBA states from each LDBA state $b_{i}\in\mathcal{B}$. At its core, our method proposes a prior distribution over these categoricals, such that appropriate shaping of the prior controls and directs the exploration in the product MDP. As the conjugate prior to categorical distributions, we adopt a Dirichlet prior $K_{i}\sim P(K_{i})=\textbf{Dir}(\alpha_{i,0},\dots,\alpha_{i,|\mathcal{B}|-1})$. The prior is informed of the LDBA structure by setting $\alpha_{i,j}=0$ if the LDBA does not allow transitions from $b_{i}$ to $b_{j}$. For the $m$ remaining non-zero Dirichlet parameters, we adopt a partially symmetric prior by setting them to a shared value $\frac{\alpha}{m}$, where $\alpha$ is a hyperparameter controlling the strength of the prior: large values of $\alpha$ induce slower convergence of the posterior distribution to the empirical transition kernel.
We remark that the choice of symmetry corresponds to the assumption that, at each step, the agent is capable of transitioning to each adjacent LDBA state with equal probability. In practice, the agent might actually take several steps in the underlying MDP in order to transition to any different LDBA state; moreover, some LDBA transitions can be substantially harder to achieve than others. Finally, for complex problems, naive exploration may not even result in observing the full set of possible transitions in a practical number of time steps. However, assuming that the agent is capable of transitioning under a max-entropy distribution allows reward signals to propagate to all states, thus resulting in informative estimates for the high-level value $\bar{V}_{K}$ and the intrinsic rewards shaping terms $R_{\text{intr}}$. For the illustrative example $T_{0}$, this is displayed in the bottom part of Figure 3.
@@
for each $i\in[0,\dots,|\mathcal{B}|]$. This update is tractable due to the choice of a conjugate Dirichlet prior for the categorical likelihood described by the transition kernel. As training progresses, a posterior distribution $P(K|D)$ can be updated with collected evidence. In practice, the updated posterior for state $b_{i}\in\mathcal{B}$ is simply $P(K_{i}|D)=\textbf{Dir}(\alpha_{i,0}+c_{i,0},\dots,\alpha_{i,|\mathcal{B}|-1}+c_{i,|\mathcal{B}|-1})$, where $c_{i,j}$ is the count of LDBA transitions from state $b_{i}$ to state $b_{j}$ observed so far. Moreover, instead of computing high-level values for a specific transition kernel $K$ as in Equation 4, we can compute the expected value $\bar{V}$ under the posterior distribution of transition kernels $P(K|D)$ and thus of MRPs:
@@
We remark that the maximization of $\bar{V}$ corresponds to the average-case MDP problem, while other procedures can be easily adapted to solve the Robust MDP (Xu & Mannor, 2010) or the Percentile Criterion MDP problems (Delage & Mannor, 2010). In practice, the expectation in Equation 7 can be estimated by sampling, with the special case of Thompson Sampling when a single sample is used. Our approach to estimating the transition kernel is ablated empirically in Appendix F; a study of the hyperparameter $\alpha$ controlling prior strength is in Appendix G.